Mr.Smith left for Buffalo for Boston; Mr. Brown using the same highway left Boston an hour later for Buffalo. These men traveled until they met. If Mr. Smith traveled at 40 mph and Mr. Brown traveled at 35 mph, find the traveling time of each man. The distance between the two cities is 460 miles.
Found 3 solutions by jorel555, advanced_Learner, MathTherapy:
Answer by jorel555(1290)

If Mr. Smith left FROM Buffalo to Boston; and vice-versa for Mr. Brown, then, let h be the number of hours before they meet. So:
40h+35h=460-1(40)
75h=420
h=5.6 hours before they meet. ☺☺☺☺

Answer by advanced_Learner(501)

mr smith
speed 40mph
distance 460
time t


460=40t
brown
speed 35
distance 460
time t+1
(35t+35)=40t
7 hours and 8 hours respectively

Answer by MathTherapy(10553)

Mr.Smith left for Buffalo for Boston; Mr. Brown using the same highway left Boston an hour later for Buffalo. These men traveled until they met. If Mr. Smith traveled at 40 mph and Mr. Brown traveled at 35 mph, find the traveling time of each man. The distance between the two cities is 460 miles.